### v1

# MultiAiTutor: Child-Friendly Educational Multilingual Speech Generation Tutor with LLMs

Abstract.
Generative speech models have demonstrated significant potential in personalizing teacher-student interactions, offering valuable real-world applications for language learning in children’s education. However, achieving high-quality, child-friendly speech generation remains challenging, particularly for low-resource languages across diverse languages and cultural contexts. In this paper, we propose MultiAiTutor, an educational multilingual generative AI tutor with child-friendly designs, leveraging LLM architecture for speech generation tailored for educational purposes. We propose to integrate age-appropriate multilingual speech generation using LLM architectures, facilitating young children’s language learning through culturally relevant image-description tasks in three low-resource languages: Singaporean-accent Mandarin, Malay, and Tamil. Experimental results from both objective metrics and subjective evaluations demonstrate the superior performance of the proposed MultiAiTutor compared to baseline methods.
Multilingual speech generations, text-to-speech synthesis, multicultural learning, AI for education.
$c$s: Applied computing Collaborative learningccs: Human-centered computing Text inputccs: Human-centered computing Auditory feedbackccs: Human-centered computing Natural language interfaces
1. Introduction
Artificial intelligence for education (AI4Edu) (Lin et al., 2023; Macina et al., 2023) has garnered increasing attention for its wide range of applications (MacLellan and Koedinger, 2022; Nye et al., 2023), including enhancing teaching and learning processes (Pokrivčáková, 2019), addressing teacher shortages through AI-based tutoring systems (Graesser et al., 2012; Chaffar and Frasson, 2004; Harley et al., 2015; Kulik and Fletcher, 2016; Nye et al., 2014), enabling dynamic online assessments (Grivokostopoulou et al., 2017), and supporting both online and offline e-learning environments (Qiao and Zhao, 2023; Zhang and Aslan, 2021). Within AI4Edu, generative speech models (Nose et al., 2007; Gao et al., 2024; Yasuda and Toda, 2023; Gao et al., 2025a; Um et al., 2020; Li et al., 2024; Wang et al., 2018) play a critical role in facilitating language acquisition, enabling children to engage effectively in listening and speaking activities across multiple languages and cultural contexts (Zhang and Aslan, 2021; Pine et al., 2025; Chen et al., 2024b).
Figure 1. The overview of the proposed MultiAiTutor for Singaporean-accented Mandarin (SG Mandarian), Malay and Tamil in multicultural contexts. Transcribed English text is provided for all three languages to aid comprehension.
Despite advances in adult speech synthesis (Lei et al., 2021; Chen et al., 2023; Liu et al., 2024; Chen et al., 2024a), child-specific speech data remains scarce (Zhang et al., 2025), making children one of the most underrepresented groups in speech generation. This scarcity arises primarily from ethical and practical challenges associated with collecting and annotating children’s speech (Zhang et al., 2025; Kulkarni et al., 2025). Furthermore, children’s speech exhibits unique acoustic and linguistic characteristics such as higher pitch, smaller vocal tracts, and evolving pronunciation patterns distinct from adult speech (Zhang et al., 2025; Kulkarni et al., 2025). Generative speech models often lack adequate child-focused training data, an issue particularly acute in low-resource languages (Gong et al., 2024).
Multilingual education is vital in culturally diverse societies like Singapore, where the national curriculum expects children to learn multiple languages—English, Singaporean-accented Mandarin, Malay, and Tamil—to preserve heritage and foster cross-cultural communication (Huzaifah et al., 2024b, a; He et al., 2024; Lovenia et al., 2024; Zhang et al., 2021). While English dominates daily use, the other three are low-resource in terms of digital tools and learning materials and remain crucial for academic assessment and cultural inclusion (Huzaifah et al., 2024b, a; He et al., 2024; Lovenia et al., 2024). However, existing educational explorations often fall short in supporting consistent and effective multilingual language learning, especially for under-resourced languages in child-centric contexts (Gong et al., 2024). The only prior work on speech generation for low-resource languages in educational contexts relies on conventional architectures such as FastSpeech (Ren et al., 2019) and FastPitch (Łańcucki, 2021), targeting languages like Onkwawenna Kentyohkwa and Kanyen’kéha in the UK and US (Kulkarni et al., 2025). In contrast, speech generation for Southeast Asian low-resource languages—with 671 million people for 8.75% of the global population (Lovenia et al., 2024)—remains largely underexplored, particularly with the use of state-of-the-art large language model (LLM) architectures over conventional neural networks.
To bridge this gap, there is a pressing need for a child-friendly, multilingual AI tutor that can support language learning in low-resource languages and culturally aware manner. In this paper, we propose MultiAiTutor, a multilingual speech generation tutor powered by LLMs, designed specifically to support culturally grounded learning experiences, age-appropriate language learning for children. Our MultiAiTutor integrates child-friendly, age-appropriate speech generation using an LLM-based neural architecture, leveraging LLM’s in-context learning to flexibly adapt to different language inputs and learner needs. This mechanism enables culturally grounded, speech-based language practice through a multicultural picture-description task in three low-resource languages, allowing children to learn anytime and anywhere.
The main contributions of this paper are: (1) we propose the first multilingual speech generation approach leveraging an LLM-based architecture; (2) we introduce an age-appropriate training strategy that advances AI4Edu by supporting child-friendly speech generation for low-resource language learning; and (3) extensive experiments show that the proposed MultiAiTutor significantly outperforms baseline models.
2. Methodology: MultiAiTutor
We propose MultiAiTutor, a child-friendly multilingual speech generation approach specifically designed to support language acquisition for low-resource languages using an AI-driven educational tutoring system. This approach enables young children to engage in age-appropriate speaking and listening activities across Tamil, Singaporean-accented Mandarin, and Malay, aligning closely with Singapore’s diverse linguistic and cultural landscape. Figure 1 illustrates the overall architecture of the MultiAiTutor approach in a multicultural setting.
2.1. Overview
As part of the learning interaction, MultiAiTutor in Figure 1 incorporates an interactive picture description task, where the AI tutor presents an image accompanied by a multilingual text and the corresponding speech, ”What do you see in the picture?” in one of the three low-resource languages. This task promotes children’s oral expression, builds vocabulary, and boosts confidence in multilingual communication through culturally relevant visuals and multi-round tutor-student conversations. The ideal student reply is illustrated in Figure 1.
Our MultiAiTutor begins with the image event captioning module, which analyzes visual inputs to generate descriptive keywords tailored for educational purposes. These keywords subsequently serve as inputs to multilingual dialogue LLMs, which generate structured responses to guide students toward replies like the example shown in Figure 1. Our multilingual speech generation, multilingual flow-matching model, and vocoder then convert these texts into child speech across all three target languages. Following this, the student verbally responds, and their speech is captured and transcribed by a multilingual speech recognition module. The resulting transcript is utilized to inform the next cycle of dialogue and speech generation by the AI tutor, ensuring continuous and adaptive engagement.
2.2. Child-friendly Generative LLMs Modeling
Motivated by the successful integration of LLMs in emotional speech synthesis (Gao et al., 2024, 2025b), we propose an innovative, child-friendly, and age-appropriate training strategy for multilingual speech generation. Our method introduces LLM architectures for the first time targeted at low-resource languages: Malay, Tamil, and Singaporean-accented Mandarin.
Unlike traditional speech generation models like Fastspeech (Kulkarni et al., 2025; Ren et al., 2019), our MultiAiTutor integrates LLM and captures aspects of children’s speech, such as higher pitch, distinctive prosodic patterns, and intonation, in a data-driven manner. This child-oriented approach allows MultiAiTutor to function effectively both as an educational tutor and as a relatable conversational friend. Leveraging the autoregressive nature of LLMs, our MultiAiTutor generates child-like multilingual speech outputs by minimizing the Kullback–Leibler (KL) divergence between the predicted and ground-truth probability distributions of multilingual speech tokens, guided by three language identifiers: Tamil, Malay, and Singaporean-accented Mandarin. This KL-based objective encourages the model to approximate natural children’s speech more closely, fostering an immersive and culturally grounded learning experience.
To illustrate this capability, Figure 1 presents representative outputs depicting culturally significant traditional festivals for each linguistic group: Deepavali for the Indian community, Chinese New Year for the Chinese community, and Hari Raya for the Malay community, where students learn culturally relevant responses—such as buying Malay clothes for Hari Raya, writing couplets for Chinese New Year, and dancing to Indian music for Deepavali. Through culturally grounded interactions, MultiAiTutor fosters an engaging environment that eases cognitive load and builds children’s confidence and language skills through child-to-child communication, demonstrating strong potential for real-world applications in elementary schools across Southeast Asia.
Figure 2. Comparison of AB Preference Tests with 95% Confidence Intervals: (1) SingaKids vs. MultiAiTutor for Malay; (2) SingaKids vs. MultiAiTutor for Tamil; and (3) CosyBase vs. MultiAiTutor for Singaporean-accented Mandarin (SG-Man). ‘NP’ denotes no preference.
3. Experiments
This section describes datasets, baselines and experimental setups.
3.1. Datasets and Baselines
We employ two recent SOTA models as baselines: SingaKids (Liu et al., 2025) with VITS (Kim et al., 2021) architecture for Malay and Tamil, and CosyBase (CosyVoice-300M) (Du et al., 2024) for Singaporean accent Mandarin. Notably, CosyBase does not support Malay or Tamil, while SingaKids does not cover Singaporean accented Mandarin for speech generation (Du et al., 2024; Liu et al., 2025). We collected, annotated, and curated multilingual text and speech data from children and adults in three languages. The Malay children’s corpus consists of 30,000 utterances from 104 Malay speakers aged 9–16. The Malay adult data comprises 23,897 utterances from 7 speakers, while the Tamil adult corpus includes 33,830 utterances from a single speaker. The Singaporean-accented Mandarin data contains 1,400 utterances from one child speaker.
3.2. Experimental Setups
The proposed MultiAiTutor integrates transformer-based LLM architectures, a pretrained flow-matching model, and a pretrained HiFi-GAN vocoder (Kong et al., 2020) for multilingual speech generation. The proposed MultiAiTutor is fine-tuned from the CosyVoice-300M (Du et al., 2024) for five epochs using dynamic batching, with the best-performing checkpoint for each language selected based on validation performance. Language identifiers are set to zh, ma, and ta to represent Singaporean-accented Mandarin, Malay, and Tamil, respectively. The remaining components of MultiAiTutor are identical to those in the SingaKids design (Liu et al., 2025). We conducted extensive listening tests, including AB preference tests (Gao et al., 2020), Mean Opinion Score (MOS) tests (Gao et al., 2019), and speech intelligibility evaluations. The tests involved 30 listeners—10 native Malay, 10 native Tamil, and 10 native Singaporean Chinese speakers—who each assessed 30 child-oriented speech samples in their respective languages.
Figure 3. Comparison of speech quality and naturalness results on MOS with 95% Confidence Intervals between SingaKids and MultiAiTutor for Tamil.
4. Results and Discussion
We examine the impact of multilingual human preference on different approaches. We also assess speech quality and speech naturalness using MOS, and analyze speech intelligence through both subjective evaluations and objective Character Error Rate (CER) measurements across multilingual settings.
4.1. Multilingual Human Preference Evaluation
To evaluate the effectiveness of MultiAiTutor, we conduct multilingual human preference evaluations using AB preference tests, where native speakers of Singaporean-accented Mandarin, Malay, and Tamil choose the better speech sample between our model and a baseline (CosyBase or SingaKids). As shown in Figure 2, 83.33 % of listeners choose MultiAiTutor over the SingaKids baseline for Malay, while only 16.00 % opt for the latter. Similarly, 77.33% of native human listeners select MultiAiTutor over CosyBase for Singaporean accented Mandarin, with only 13.33 % selecting the baseline. For Tamil, 91.33 % of participants select MultiAiTutor compared to just 7.34 % for SingaKids. These findings indicate that MultiAiTutor consistently outperforms baseline models across linguistically and culturally diverse settings, highlighting the effectiveness of our age-appropriate, multilingual speech generation design powered by LLMs for educational applications.
4.2. Speech Quality and Naturalness
To evaluate speech quality and naturalness, we conduct separate MOS tests across the three target languages, asking listeners to rate the samples from 1 (bad) to 5 (excellent).
4.2.1. Generative Speech Evaluation for Tamil Language
Figure 3 presents the evaluation results for generated Tamil speech. The proposed MultiAiTutor achieves significantly higher scores in both speech quality (4.05) and naturalness (3.72) compared to the Singakids baseline, which scores 2.51 and 2.19, respectively. These results demonstrate the effectiveness of MultiAiTutor in capturing Tamil linguistic and cultural characteristics of Tamil through a generative speech model based on LLM neural architectures. This advancement underscores the model’s potential to support culturally aware speech generation and promote research for the low-resource Tamil language within the speech community.
Figure 4. Comparison of speech quality and naturalness results on MOS Scores with 95% Confidence Intervals between CosyBase and MultiAiTutor for Singaporean-accented Mandarin (SG-Man).
4.2.2. Generative Speech Evaluation for Singaporean-Accented Mandarin
Figure 4 presents the subjective evaluation results for synthesized Singaporean-accented Mandarin. The proposed MultiAiTutor achieves higher scores in both speech quality (3.97 vs. 3.05) and naturalness (4.03 vs. 3.85) compared to the CosyBase baseline. These results demonstrate the effectiveness of our age-appropriate generative speech model adaptation mechanism in producing child-friendly speech suitable for educational applications in the Singaporean-accented Mandarin context.
4.2.3. Generative Speech Evaluation for Malay Language
Figure 5 presents the evaluation results for generative Malay speech. The proposed MultiAiTutor achieves higher scores in both speech quality (3.61 vs. 2.62) and naturalness (3.63 vs. 2.46) compared to the SingaKids. These findings reflect the effectiveness of MultiAiTutor in generating culturally appropriate Malay speech, extending its applicability beyond Tamil and Mandarin to the low-resource Malay language. This advancement underscores the potential of the approach to contribute to culturally inclusive speech generation research for underrepresented languages.
4.3. Multilingual Speech Intelligence
To comprehensively assess the pronunciation accuracy of the proposed approach, we perform both objective and subjective evaluations across the three low-resource languages, as presented in Table 1. For objective evaluations, we utilize automatic speech recognition (ASR) for each language: the Tamil and Malay ASR models from (Liu et al., 2025), and the Whisper model [^1] https://huggingface.co/openai/whisper-large-v3 for Mandarin. The synthesized speech samples are transcribed using these models, and CER is calculated by comparing the transcriptions to the ground-truth text. A lower CER reflects higher pronunciation accuracy.
However, in low-resource scenarios, ASR systems often exhibit limited robustness and may not fully capture the nuances of pronunciation accuracy, particularly in children’s speech. To address these limitations, we complement the objective evaluation with human assessments. For the subjective evaluation, native speakers are asked to rate the pronunciation accuracy—referred to as human speech intelligence—on a scale from 0 (completely inaccurate) to 100 (fully accurate). These human ratings offer more culturally and linguistically grounded insights, reflecting alignment with native pronunciation norms across the respective languages.
Table 1 shows that the proposed MultiAiTutor achieves consistently better performance than the baselines across all three languages, as measured by both objective evaluations using CER and subjective evaluations by human listeners. In terms of CER, MultiAiTutor substantially reduces recognition errors compared to the baselines across three languages. These improvements reflect a higher degree of pronunciation clarity in the generated speech, making it more intelligible to ASR systems. Notably, the improvements for Tamil (from 55.08 to 83.29) and Malay (from 61.25 to 76.40) in human-rated pronunciation accuracy are substantial, highlighting the effectiveness of the proposed approach in low-resource language settings. These results confirm the capability of MultiAiTutor to advance multilingual speech generation through the integration of LLM-based modeling.
Figure 5. Comparison of Speech Quality and Naturalness Results Based on MOS Scores with 95% Confidence Intervals between SingaKids and the proposed MultiAiTutor for Malay.
Table 1. Comparison of speech intelligence results from subjective evaluation and objective evaluation by CER in Singaporean-accent Mandarin, Malay and Tamil.
Singaporean-accent Mandarin
Speech Intelligence CosyBase MultiAiTutor
Objective Evaluation: CER (%) 10.70 4.00
Subjective Evaluation: Human 74.37 83.40
Tamil
Speech Intelligence SingaKids MultiAiTutor
Objective Evaluation: CER (%) 9.60 2.10
Subjective Evaluation: Human 55.08 83.29
Malay
Speech Intelligence SingaKids MultiAiTutor
Objective Evaluation: CER (%) 3.40 2.30
Subjective Evaluation: Human 61.25 76.40
5. Conclusion
We propose a multilingual generative AI tutor, MultiAiTutor, designed specifically for child-friendly speech generation and educational purposes. Our approach leverages advanced LLM architectures and culturally relevant, age-appropriate training strategies to effectively support language learning in low-resource languages such as Malay, Tamil, and Singaporean-accented Mandarin. The proposed system fosters a supportive and engaging learning environment, reducing cognitive load and enhancing learner participation. Synthesized multilingual speech samples from different models are available for verification at the link  [^2] https://xiaoxue1117.github.io/icmi2025demo/.